# Large Language Models for Next Point-of-Interest Recommendation

### v1

@@ 1. Introduction
To exploit such contextual information in LBSN data, there are some substantial challenges: (i) How to extract the contextual information from the raw data? And (ii) How to connect contextual information with commonsense knowledge to effectively benefit next POI recommendation? Here, we consider contextual information as time, POI category, and geo-coordinates. And we define commonsense knowledge in the context of the next POI recommendation as the capability to understand the semantics of contextual information without additional data, and to connect certain joint patterns of contextual information with behaviors in the real world. Existing next POI recommendation methods (Lim et al., 2020; Zhang et al., 2022a; Yang et al., 2022; Yan et al., 2023a) have two important limitations when dealing with contextual information: (i) Due to their numerical nature, they have to transform heterogeneous LBSN data into numbers. For example, POI categories are often encoded from text into IDs. This transformation can result in a loss of inherent meaning associated with contextual information. (ii) They rely exclusively on statistics and human designs to understand contextual information and lack an understanding of the semantic concepts provided by the contextual information.
@@
More specifically, to address challenge (i), we transform the next POI task into a question-answering task. We convert text-formed raw data into prompts constructed by blocks of sentences. Each block serves as a different module, and each sentence in it contains the necessary information for that module in its original format. Therefore, all heterogeneous data can be fed into a single model with tokenization that still keeps the contextual information. As illustrated in Figure 1, unlike traditional numerical methods, where data needs to be transformed and fed into different embedding layers, our method allows the data to be directly used in its original format. We also propose a notion of trajectory similarity based on prompts, which is used for the cold-start problem.
(a) Numeric paradigm.
@@
Figure 1. Comparison of two paradigms for the next POI task: (a) a typical numerical paradigm and (b) the proposed language model-based paradigm.
@@
Table 1. POI categories in a real-world dataset summarized by ChatGPT.
Category Subcategories/Category Names
@@
    We propose a framework to use pretrained large language models for the next POI recommendation task, which brings commonsense knowledge for making use of the rich contextual information in the data. To the best of our knowledge, we are the first to fine-tune language models on standard-sized datasets to exploit commonsense knowledge for the next POI recommendation task.
@@ 2.1. Next POI Recommendation
Subsequent work has begun to apply RNN-based models with the rise of deep learning. Kong and Wu (2018) proposed HST-LSTM, where they added spatial-temporal factors into LSTM gates to guide the learning and further employed a hierarchical extension to model the periodicity of the visit sequence. LSTPM (Sun et al., 2020) employed three LSTM modules, utilizing non-local neural operations and a short-term preference modeling module using geo-dilated LSTM. PLSPL (Wu et al., 2020) combined an embedding layer with the attention mechanism to learn long-term preferences and leveraged two LSTM models to model short-term preferences at both the location and category levels. STAN (Luo et al., 2021) uses a multimodal embedding layer to learn the representation of user, location, time, and spatial-temporal effect, with a bi-layer attention architecture to learn the explicit spatial-temporal relevance within the trajectories. CFPRec (Zhang et al., 2022b) focused on the multi-step future plan of users by adopting an attention mechanism to extract future references from past and current preference encoders that are transformer and LSTM encoders. These sequence-based models often become confined to a local view and also suffer from short trajectories and the cold-start problem where inactive users have limited data. Our method deploys key-query similarity, which allows us to combine information from different users, alleviating the cold-start problem.
Graph-based models. More recently, graph-based methods have been incorporated to address the limitations of sequence-based models. STP-UDGAT (Lim et al., 2020) were the first to use a graph attention network (Veličković et al., 2018), enabling users to selectively learn from others in a global view. Zhang et al. (2022a) proposed a hierarchical multi-task graph recurrent network (HMT-GRN) to learn user-POI and user-region distribution, employing a GRN to replace the LSTM unit to learn both sequential and global spatial-temporal relationships between POIs. DRGN (Wang et al., 2022) investigated the intrinsic characteristics of POIs by learning disentangled representation from both distance-based and transition-based relation graphs through a GCN layer. GETNEXT (Yang et al., 2022) addressed the cold start problem by exploiting collaborative signals from other users and proposing a global trajectory flow map and a novel graph-enhanced transformer model. STHGCN (Yan et al., 2023a) alleviated cold-start issues by constructing a hypergraph to capture higher-order information, including user trajectories and collaborative relations. Although graph-based models handle the cold-start problem, they are not capable of combining contextual information with commonsense knowledge. Our method avoids contextual information loss by trajectory prompting, and the LLMs that we use contains commonsense knowledge to understand contextual information.
@@ 4.2. Key-Query Pair Similarity
Figure 4. Similarity computation for each pair of key and query. Each pair of key and query prompts are fed into a LLAMA2 encoder separately. We use the last hidden layer embeddings to compute their cosine similarity.
@@ 5.1.1. Datasets
We conduct experiments on three public datasets: Foursquare-NYC, Foursquare-TKY (Yang et al., 2014), and Gowala-CA (Cho et al., 2011). The first two datasets, collected over 11 months, comprise data from New York City and Tokyo, sourced from Foursquare. The Gowala-CA dataset, from the Gowalla platform, covers a broader geographical area and time period, encompassing California and Nevada. We utilize data that has been preprocessed as per the methods detailed by Yan et al. (2023b). The data is preprocessed as follows: (i) Filter out Points of Interest (POIs) with fewer than 10 visit records in history; (ii) Exclude users with fewer than 10 visit records in history; (iii) Divide user check-in records into several trajectories with 24-hour intervals, excluding trajectories that contain only one check-in record. The check-in records are also sorted chronologically: the first 80% are used for the training set, the middle 10% are defined as the validation set, and the last 10% are defined as the test set. Note that the validation and test set has to contain all users and POIs that appear in the training set. The unseen users and POIs would be removed from the validation and test set.
@@
  • STHGCN (Yan et al., 2023b): Constructing a hypergraph to capture inter and intra-user relations, STHGCN proposes a hypergraph transformer and solves the cold-start problem.
@@ 5.1.3. Our Models
In our experiments we consider three versions of what we call “our model”: (i) LLAMA2-7b: Our model, using prompts only the current trajectory block without the historical trajectory block, where we use Llama-2-7b-longlora-32k (Chen et al., 2023; Touvron et al., 2023) as our base LLM. (ii) LLAMA2-7b*: A variation on LLAMA2-7b where we use prompts with the historical trajectory block without applying key-query similarity, where only the historical trajectories from the current users are considered. (iii) LLAMA2-7b**: A second variation on LLAMA2-7b where we use prompts with the historical trajectory block combined with key-query similarity, where historical trajectories from both the current users and other users are considered. Below, unless stated otherwise, when we say “our model,” we refer to the LLAMA2-7b** variant.
@@ 5.1.4. Evaluation Metrics
For evaluation we regard the next POI recommendation task as a top-1 recommendation problem. The scenario we have in mind is one with “extreme position bias” (Fairstein et al., 2022), where only a small amount of information can be presented to the user during a single interaction. For example, a user who is on a business trip wishes to explore the new city before the meeting. In such a scenario, the user does not have the leisure to review multiple options. This is a scenario that is usually considered in next POI recommendation, often as the primary scenario (see, e.g., Yan et al., 2023b; Yang et al., 2022). Given this choice of scenario, our approach to the task as a type of question-answering problem is a natural fit. We prioritize the delivery of the most pertinent and contextually suitable recommendation, mirroring the objective of providing a single, correct answer to a user’s query.
@@ 5.1.5. Implementation Details
LLAMA2-7b $\times$ $\times$ 0.2356 0.1517 0.1016
LLAMA2-7b* $\times$ 0.3171 0.2836 0.1683
LLAMA2-7b** 0.3372 0.3035 0.2065
@@ 5.3.1. User Cold-start Analysis
Inactive Ours 0.3417 0.3478 0.2132
Normal Ours 0.3841 0.3516 0.2057
Very active Ours 0.3088 0.2727 0.1920
@@ 5.3.2. Trajectory Length Analysis
Short Ours 0.3364 0.2876 0.1955
Middle Ours 0.3350 0.3013 0.1998
Long Ours 0.3271 0.3083 0.2037
@@ 5.3.3. Number of Historical Data Variants
Table 6. Analysis on NYC, TKY, and CA dataset for LLAMA2-7b** trained on prompts with different numbers of historical trajectories.
@@ 5.3.4. Generalization to Unseen Data Analysis
Table 7. The models are LLAMA2-7b** trained only on one of the NYC, TKY, and CA datasets and evaluated on the rest.
@@ 5.3.5. Contextual Information Analysis
Table 8 shows the overall results for the models trained in NYC, TKY, and CA and tested with different prompts. We can see that the model performance drops by a small margin when we remove the contextual information from POI category names in NYC. For TKY and CA, the model performance decreases by 6.4% and 6.2%, respectively.
Table 8. The results for LLAMA2-7b** tested with different prompts.
@@
Table 9. The results for LLAMA2-7b** tested with different prompts in terms of users with different levels of activity.
@@ 5.3.6. Effect of Different Components
Table 10. Ablation study results for LLAMA2-7b** over three datasets.
@@ 6. Conclusion and Future Work
In this paper, we propose a framework to deploy large language models for the next point-of-interest recommendation task, which is the first to utilize models with commonsense knowledge for the task. We developed trajectory prompting to transform the task into a question-answering. We also introduce key-query similarity to alleviate the cold-start problem.
@@
Acknowledgement
This research was conducted by the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005), and funded by the Australian Government through the Australian Research Council. In addition, this research was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.

### v2

@@ 1. Introduction
To exploit such contextual information in LBSN data, there are some substantial challenges: (i) How to extract the contextual information from the raw data? And (ii) How to connect contextual information with commonsense knowledge to effectively benefit next POI recommendation? Here, we consider contextual information as time, POI category, and geo-coordinates. And we define commonsense knowledge in the context of the next POI recommendation as the capability to understand the semantics of contextual information without additional data, and to connect certain joint patterns of contextual information with behaviors in the real world. Existing next POI recommendation methods (Lim et al., 2020; Zhang et al., 2022; Yang et al., 2022; Yan et al., 2023) have two important limitations when dealing with contextual information: (i) Due to their numerical nature, they have to transform heterogeneous LBSN data into numbers. For example, POI categories are often encoded from text into IDs. This transformation can result in a loss of inherent meaning associated with contextual information. (ii) They rely exclusively on statistics and human designs to understand contextual information and lack an understanding of the semantic concepts provided by the contextual information.
@@
More specifically, to address challenge (i), we transform the next POI task into a question-answering task. We convert text-formed raw data into prompts constructed by blocks of sentences. Each block serves as a different module, and each sentence in it contains the necessary information for that module in its original format. Therefore, all heterogeneous data can be fed into a single model with tokenization that still keeps the contextual information. As illustrated in Figure 1, unlike existing methods that we categorize as feature transformation paradigms, where data needs to be transformed and fed into different embedding layers, our method allows the data to be directly used in its original format. We also propose a notion of trajectory similarity based on prompts, which is used for the cold-start problem.
(a) Feature transformation paradigm.
@@
Figure 1. Comparison of two paradigms for the next POI task: (a) a typical feature transformation paradigm and (b) the proposed language model-based paradigm.
@@
Table 1. POI categories in a real-world dataset summarized by ChatGPT. The left column is summarized by ChatGPT. We first feed the raw csv file to ChatGPT and ask it to find the unique POI category names. Then we use the prompt ‘Can you list the category names that have intersections by their context?’ to generate the categories that are summarized from the right column, which represents the subcategories from the raw data.
Category (GPT) Subcategories/Category Names (Raw Data)
@@
    We propose LLM4POI, a framework to use pretrained large language models for the next POI recommendation task, which brings commonsense knowledge for making use of the rich contextual information in the data. To the best of our knowledge, we are the first to fine-tune language models on standard-sized datasets to exploit commonsense knowledge for the next POI recommendation task.
@@ 2.1. Next POI Recommendation
Subsequent work has begun to apply RNN-based models with the rise of deep learning. Kong and Wu (2018) proposed HST-LSTM, where they added spatial-temporal factors into LSTM gates to guide the learning and further employed a hierarchical extension to model the periodicity of the visit sequence. LSTPM (Sun et al., 2020) employed three LSTM modules, utilizing non-local neural operations and a short-term preference modeling module using geo-dilated LSTM. PLSPL (Wu et al., 2020) combined an embedding layer with the attention mechanism to learn long-term preferences and leveraged two LSTM models to model short-term preferences at both the location and category levels. STAN (Luo et al., 2021) uses a multimodal embedding layer to learn the representation of user, location, time, and spatial-temporal effect, with a bi-layer attention architecture to learn the explicit spatial-temporal relevance within the trajectories. CFPRec (Zhang et al., 2022) focused on the multi-step future plan of users by adopting an attention mechanism to extract future references from past and current preference encoders that are transformer and LSTM encoders. These sequence-based models often become confined to a local view and also suffer from short trajectories and the cold-start problem where inactive users have limited data. Our method deploys key-query similarity, which allows us to combine information from different users, alleviating the cold-start problem.
Graph-based models. More recently, graph-based methods have been incorporated to address the limitations of sequence-based models. STP-UDGAT (Lim et al., 2020) were the first to use a graph attention network (Veličković et al., 2018), enabling users to selectively learn from others in a global view. Zhang et al. (2022) proposed a hierarchical multi-task graph recurrent network (HMT-GRN) to learn user-POI and user-region distribution, employing a GRN to replace the LSTM unit to learn both sequential and global spatial-temporal relationships between POIs. DRGN (Wang et al., 2022) investigated the intrinsic characteristics of POIs by learning disentangled representation from both distance-based and transition-based relation graphs through a GCN layer. GETNEXT (Yang et al., 2022) addressed the cold start problem by exploiting collaborative signals from other users and proposing a global trajectory flow map and a novel graph-enhanced transformer model. STHGCN (Yan et al., 2023) alleviated cold-start issues by constructing a hypergraph to capture higher-order information, including user trajectories and collaborative relations. Although graph-based models handle the cold-start problem, they are not capable of combining contextual information with commonsense knowledge. Our method avoids contextual information loss by trajectory prompting, and the LLMs that we use contains commonsense knowledge to understand contextual information.
@@ 4.2. Key-Query Pair Similarity
Figure 4. Similarity computation for each pair of key and query. Each pair of key and query prompts is fed into a pre-trained LLAMA2 separately. We use the last hidden layer embeddings to compute their cosine similarity.
@@
Figure 5. For data in the training set, the ¡question¿ part in the prompt is fed into the pre-trained LLM, and the ¡answer¿ part is used to guide the fine-tuning.
@@ 5.1.1. Datasets
We conduct experiments on three public datasets: Foursquare-NYC, Foursquare-TKY (Yang et al., 2014), and Gowala-CA (Cho et al., 2011). The first two datasets, collected over 11 months, comprise data from New York City and Tokyo, sourced from Foursquare. The Gowala-CA dataset, from the Gowalla platform, covers a broader geographical area and time period, encompassing California and Nevada. We utilize data that has been preprocessed as per the methods detailed by Yan et al. (2023). The data is preprocessed as follows: (i) Filter out Points of Interest (POIs) with fewer than 10 visit records in history; (ii) Exclude users with fewer than 10 visit records in history; (iii) Divide user check-in records into several trajectories with 24-hour intervals, excluding trajectories that contain only one check-in record. The check-in records are also sorted chronologically: the first 80% are used for the training set, the middle 10% are defined as the validation set, and the last 10% are defined as the test set. Note that the validation and test set has to contain all users and POIs that appear in the training set. The unseen users and POIs would be removed from the validation and test set.
@@
  • STHGCN (Yan et al., 2023): Constructing a hypergraph to capture inter and intra-user relations, STHGCN proposes a hypergraph transformer and solves the cold-start problem.
@@ 5.1.3. Our Models
In our experiments we consider three versions of what we call “our model”: (i) LLM4POI*: Our model, using prompts only the current trajectory block without the historical trajectory block, where we use Llama-2-7b-longlora-32k (Chen et al., 2023; Touvron et al., 2023) as our base LLM. (ii) LLM4POI**: A variation on LLAMA2-7b where we use prompts with the historical trajectory block without applying key-query similarity, where only the historical trajectories from the current users are considered. (iii) LLM4POI: A second variation on LLAMA2-7b where we use prompts with the historical trajectory block combined with key-query similarity, where historical trajectories from both the current users and other users are considered. Below, unless stated otherwise, when we say “our model,” we refer to the LLM4POI variant.
@@ 5.1.4. Evaluation Metrics
For evaluation we regard the next POI recommendation task as a top-1 recommendation problem. The scenario we have in mind is one with “extreme position bias” (Fairstein et al., 2022), where only a small amount of information can be presented to the user during a single interaction. For example, a user who is on a business trip wishes to explore the new city before the meeting. In such a scenario, the user does not have the leisure to review multiple options. This is a scenario that is usually considered in next POI recommendation, often as the primary scenario (see, e.g., Yan et al., 2023; Yang et al., 2022). Given this choice of scenario, our approach to the task as a type of question-answering problem is a natural fit. We prioritize the delivery of the most pertinent and contextually suitable recommendation, mirroring the objective of providing a single, correct answer to a user’s query.
@@ 5.1.5. Implementation Details
LLM4POI* $\times$ $\times$ 0.2356 0.1517 0.1016
LLM4POI** $\times$ 0.3171 0.2836 0.1683
LLM4POI 0.3372 0.3035 0.2065
@@ 5.3.1. User Cold-start Analysis
Inactive LLM4POI 0.3417 0.3478 0.2132
Normal LLM4POI 0.3841 0.3516 0.2057
Very active LLM4POI 0.3088 0.2727 0.1920
@@ 5.3.2. Trajectory Length Analysis
Short LLM4POI 0.3364 0.2876 0.1955
Middle LLM4POI 0.3350 0.3013 0.1998
Long LLM4POI 0.3271 0.3083 0.2037
@@ 5.3.3. Number of Historical Data Variants
Table 6. Analysis on NYC, TKY, and CA dataset for LLM4POI trained on prompts with different numbers of historical trajectories.
@@ 5.3.4. Generalization to Unseen Data Analysis
Table 7. The models are LLM4POI trained only on one of the NYC, TKY, and CA datasets and evaluated on the rest.
@@ 5.3.5. Contextual Information Analysis
Table 8 shows the overall results for the models trained on NYC, TKY, and CA and tested with different prompts. Model performance drops by a small margin when we remove the contextual information from POI category names in NYC. For TKY and CA, the model performance decreases by 6.4% and 6.2%, respectively.
Table 8. The results for LLM4POI tested with different prompts.
@@
Table 9. The results for LLM4POI tested with different prompts in terms of users with different levels of activity.
@@ 5.3.6. Effect of Different Components
Table 10. Ablation study results for LLM4POI over three datasets.
@@ 6. Conclusion and Future Work
In this paper, we propose LLM4POI, a framework to deploy large language models for the next point-of-interest recommendation task, which is the first to utilize models with commonsense knowledge for the task. We developed trajectory prompting to transform the task into a question-answering. We also introduce key-query similarity to alleviate the cold-start problem.
@@
Acknowledgements
This research was conducted by the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005), and funded by the Australian Government through the Australian Research Council. In addition, this research was undertaken with the assistance of resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government. Peibo Li is supported by Google Conference Scholarships (APAC). Maarten de Rijke was partially supported by the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006.
All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.